# Benchmarking and Improving Large Vision-Language Models for Fundamental Visual Graph Understanding and Reasoning

### v1

@@ Abstract
Large Vision-Language Models (LVLMs) have demonstrated remarkable performance across diverse tasks. Despite great success, recent studies show that LVLMs encounter substantial limitations when engaging with visual graphs. To study the reason behind these limitations, we propose VGCure, a comprehensive benchmark covering 22 tasks for examining the fundamental graph understanding and reasoning capacities of LVLMs. Extensive evaluations conducted on 14 LVLMs reveal that LVLMs are weak in basic graph understanding and reasoning tasks, particularly those concerning relational or structurally complex information. Based on this observation, we propose a structure-aware fine-tuning framework to enhance LVLMs with structure learning abilities through 3 self-supervised learning tasks. Experiments validate the effectiveness of our method in improving LVLMs’ zero-shot performance on fundamental graph learning tasks, as well as enhancing the robustness of LVLMs against complex visual graphs [^1] The code and datasets are available at https://github.com/AAAndy-Zhu/VGCure.
@@
Yingjie Zhu1,2, Xuefeng Bai1, Kehai Chen1, Yang Xiang2, Min Zhang1 1Institute of Computing and Intelligence, Harbin Institute of Technology, Shenzhen, China 2 Peng Cheng Laboratory, Shenzhen, China zhuyj@stu.hit.edu.cn, {chenkehai,baixuefeng,zhangmin2021}@hit.edu.cn xiangy@pcl.ac.cn
@@ 1 Introduction
Graphs serve as a fundamental data structure across a wide range of domains, including social network analysis Schweimer et al. (2022), recommendation systems Zhang et al. (2023), knowledge graphs Zhang et al. (2024b), chemistry Cao et al. (2024) and biomedical molecules Liu et al. (2023). Existing methods have achieved great success in enhancing understanding and reasoning abilities in graph-based tasks Kim et al. (2023); Chen et al. (2024). However, these approaches typically focus on specific graph types or tasks, posing challenges in designing versatile systems that are suitable for various tasks and graphs across diverse domains.
@@
Recently, Large Vision-Language Models (LVLMs) have exhibited outstanding performance across a diverse range of downstream tasks by unifying various inputs in the form of images and processing them with human-like understanding and reasoning abilities Zhu et al. (2024); Zheng et al. (2024). This triggers an increasing interest in employing LVLMs for graph learning problems, given the vision modality offers a natural and intuitive way for comprehending structural information and facilitating general graph reasoning Poklukar et al. (2022). Despite significant advancements, recent studies Wei et al. (2024); Li et al. (2024d); Ai et al. (2024) reveal that LVLMs face significant challenges in addressing graph-based learning problems. Notably, LVLMs achieve less than 15% accuracy on mathematical graph reasoning tasks, markedly below their performance in image and text reasoning Li et al. (2024d). Therefore, a natural challenge arises: why do LVLMs fail in graph learning, and how to enhance LVLMs to process graphs like professionals?
@@
To address this challenge, we begin by identifying the research gap in existing evaluations that limit our understanding of LVLMs. Current work primarily assesses LVLMs on graph theory problems or multi-hop reasoning tasks Wei et al. (2024); Ai et al. (2024), which are inherently complex and necessitate a combination of diverse cognitive abilities. Thus it remains uncertain whether LVLMs have strong fundamental understanding and reasoning capabilities in processing visual graphs—such as recognizing the basic components of a graph and making basic logical inferences based on graph data—which is crucial for determining whether the limitations of LVLMs originate from a lack of fundamental abilities or other higher-order capabilities.
@@
To this end, we present the Vision Graph Comprehensive Understanding and REasoning benchmark, VGCure, to thoroughly evaluate the fundamental understanding and reasoning capabilities of LVLMs concerning visual graphs. As shown in Figure 1, VGCure evaluates the fundamental capabilities of LVLMs across a comprehensive set of tasks, including 9 graph understanding tasks and 13 graph reasoning tasks. Moreover, VGCure encompasses 10 anonymized graph structures sourced from both synthetic and real-world sources, offering a robust testbed for assessing LVLMs’ proficiency in handling diverse graphs. We conduct extensive experiments involving 14 representative LVLMs using VGCure. Experiments show that 1) LVLMs exhibit subpar performance in most tasks, especially in capturing relational information, indicating their weak fundamental understanding and reasoning skills on visual graphs; 2) LVLMs are frequently confused when confronted with structurally complex visual graphs.
@@
Motivated by the above observations, we further introduce a novel self-supervised, structure-aware fine-tuning framework MCDGraph to enhance the structure learning capabilities of LVLMs. In particular, MCDGraph comprises three tasks: 1) masked graph infilling, which guides LVLMs to infer missing local structural information; 2) contrastive graph discrimination, which guides LVLMs to distinguish structural differences between graphs; 3) graph description, which teaches LVLMs to summarize the global information inherent in graph structures. Experiments show that our self-supervised framework considerably enhances the graph understanding and reasoning abilities of two representative LVLMs, particularly in their performance on edge-related tasks. Further analysis demonstrates that our method improves LVLMs’ robustness in handling more structurally complex visual graphs. The contributions of this work can be summarized as follows:
@@
  • Through extensive experiments on 14 LVLMs, we reveal LVLMs’ limitations in fundamental graph understanding and reasoning, especially for tasks concerning relational or structural information. Besides, LVLMs’ performance declines as graph complexity increases.
  • We propose a self-supervised framework to enhance LVLMs’ ability in capturing the structural information in visual graphs. Experiments validate the effectiveness of our approach across a wide range of tasks.
@@
Table 2: Examples for each task in VGCure. These samples all correspond to the graph shown in Figure LABEL:fig:example_graphs(d).
@@ 2 The VGCure Benchmark
To evaluate LVLMs’ capabilities in fundamental graph understanding and reasoning, we introduce VGCure, a large-scale multimodal graph benchmark with 22 challenging tasks in two categories. VGCure features 10 graph types, both synthetic and real-world, to assess LVLMs’ effectiveness with diverse graphs. It anonymizes the graphs to minimize the impact of pre-existing LLM knowledge on core reasoning abilities, promoting knowledge-free reasoning (Hu et al., 2024). Table 1 offers a comparison between VGCure and three recent benchmarks. It is evident that VGCure excels in fundamental graph understanding and reasoning capabilities. Furthermore, VGCure offers a comprehensive evaluation through more varied graph types, tasks, and test samples.
@@ 2.1 Graph Structure Generation
We begin by collecting a wide variety of graph structures that we will then use to generate visual graphs and challenging tasks. Following Fatemi et al. (2024), we first leverage NetworkX library (Hagberg et al., 2008) for generating a diverse set of random synthetic structures, including Erdős-Rényi (ER) graphs (ERDdS and R&wi, 1959), scale-free networks (SFN) (Barabási and Albert, 1999), Barabási–Albert (BA) model (Albert and Barabási, 2002), stochastic block model (SBM) (Holland et al., 1983) and star graphs. In addition, we extract anonymized structures from real-world graphs in GraphArena benchmark (Tang et al., 2024), including DBLP, Social Network, DBpedia, Openflights and PubChemQC. All the entity and relation names in each graph are replaced with generic names to eliminate the impact of the model’s internal knowledge on reasoning. After initializing the graph structure, we use the Graphviz library (Ellson et al., 2002) to generate concise directed and undirected visual graphs.
@@ 2.2 Tasks Design
To thoroughly assess the abilities of LVLMs in fundamental graph structure understanding and reasoning, the proposed VGCure encompasses the following categories of tasks. Table 2 presents examples for each task.
Category 1: Graph Understanding. The graph understanding tasks involve analyzing and extracting structural, relational, and property-based information from the visual graph, which aims to gain insights into the composition and topology of the graph, including its nodes, edges, connectivity, and the relationships among its components.
@@
  • Degree Query (DQ): Calculate the degree of the given node, representing the number of edges connected to it.
@@
Category 2: Graph Reasoning. The reasoning tasks focus on exploring the knowledge-free reasoning ability of LVLMs on visual graphs. To differentiate from graph understanding tasks, we designed a series of 2-hop reasoning tasks.
@@
  • Conjunctive Relation Query (CR): Retrieve the entities that simultaneously satisfied two independent relationship constraints with two distinct entities.
@@
For each task, we construct one QA sample and two fact checking (FC) samples (with labels of True and False, respectively) automatically based on the template, except for CN and DC which have only fact checking samples due to the strong similarity between the two samples. The total number of final samples is 223,646. More examples and details about VGCure benchmark can be found in Appendix A.
@@ 3.1 Experimental Setup
We conduct evaluation on 13 open-source LVLMs, including InternLM-XComposer2.5-7B (Zhang et al., 2024a), InternVL2-8B, Llama3.2-11B-Vision-Instruct, LLaVA1.5-7B (Liu et al., 2024), LLaVA1.5-13B (Liu et al., 2024), LLaVA-NeXT-7B (Li et al., 2024b), LLaVA-OneVision-7B (Li et al., 2024a), MiniGPT-v2 (Chen et al., 2023), Monkey (Li et al., 2024e), mPLUG-Owl3-7B (Ye et al., 2024), Qwen-VL (Bai et al., 2023), Qwen2-VL-7B-Instruct (Wang et al., 2024) and SPHINX (Lin et al., 2023). Meanwhile, we also evaluate the performance of the GPT-4o-mini, which is a strong closed-source LVLM [^2] We excluded GPT-4o as a baseline due to its high cost.. Due to the high cost of GPT-4o-mini, we randomly select 50 graphs from each graph structure for testing. For all methods, the zero-shot setting is adopted during evaluation. More details can be found in Appendix B.
@@ 3.2 Main Result
Graph Understanding Table 3 and Table 4 present the evaluation results in the forms of question answering (QA) and fact checking (FC), respectively. We report the averaged performance across various graph structures. In general, most LVLMs struggle to precisely understand the structural and relational information in visual graphs. In specific, (I) Among the graph understanding tasks, Node Number Query (NNu) and Directedness Check (DC) are the easiest for most LVLMs. This indicates that most LVLMs are able to accurately recognize the number of specific elements in the visual graph. (II) All LVLMs struggle with Edge Number Query (EN) and Neighbor Query (NQ) tasks, with the highest performance ratings remaining 16.28% accuracy and 30.81% F1 score, respectively. This indicates that current LVLMs are weak in understanding relational and structural information. (III) Even with a similar number of parameters, the graph understanding abilities of open-source LVLMs vary significantly, with Qwen2-VL and InternVL2 showing better performance in both QA and FC samples. (IV) For the same task, LVLMs perform differently on QA and FC samples, likely due to different ways in reasoning and understanding required by each task (Thorne et al., 2018). (V) The closed-source LVLM, GPT-4o-mini, offers no significant advantages and even underperforms open-source LVLMs, especially on FC tasks.
Graph Reasoning Based on graph reasoning results in Tables 3 and 4, it can be observed that, (I) Compared to graph understanding, the graph reasoning tasks are more challenging and the overall performance of LVLMs is worse on both QA and FC samples. This may be a knock-on effect due to deficiencies in visual graph understanding. (II) Among all open-source LVLMs, InternVL2 and Llama3.2-Vision achieve better performance and LLaVA1.5-7B perform the worst on graph reasoning tasks. (III) All the LVLMs perform poorly on Nested Relation Query (NR) for both QA and FC samples, which is similar to the observation in the graph understanding task. This suggests that LVLMs are deficient in recognizing edges and understanding structural information within visual graphs. (IV) On the QA samples, the performance of different LVLMs on Connective Path Query (CP) varies widely. SPHINX, Monkey and Qwen-VL demonstrate almost no ability to recognize paths between two specific nodes in the visual graph. (V) Similarly, GPT-4o-mini underperforms in most tasks compared to open-source LVLMs, except for the Connective Path Query (CP) task.
Figure 2: Model performance (F1/Acc) on QA samples across various graph structures and tasks , where OF, PCQC and SN denotes Openflights, PubChemQC and Social network, respectively.
@@ 3.3 Impact of Structures
Inspired by Fatemi et al. (2024), we investigate whether graph structure influences LVLMs’ ability to understand and reason over visual graphs. Figure 2 compares the performance of the top five LVLMs on QA samples across various graph structures. It is evident that graph structure significantly affects the understanding and reasoning performance of LVLMs on most tasks. Specifically, all LVLMs obtain impressive results on PCQC and weaker performance on BA, possibly due to the simpler structure of PCQC with fewer nodes and edges (averaging 5.45 nodes and 4.76 edges), while BA contains the largest number of edges in VGCure. Furthermore, on different tasks, the performance of LVLMs is affected differently by the graph structure. Edge Number Query (EN) and Connective Path Query (CP) show larger performance variations across graph structures, whereas Node Number Query (NNu) and Nested Node Query (NN) show smaller differences. As presented in Figure 17, the overall trend of FC is similar to that of QA.
@@ 3.4 Impact of Complexity
We further discuss the impact of graph complexity on the LVLMs’ ability to understand and reason over the visual graph considering the number of edges, number of nodes, and average degree of the graph. Figure LABEL:fig:complexity_results shows the performance comparison of five LVLMs on QA samples across representative tasks and complexity levels. It can be observed that as graph complexity increases, the performance of LVLMs declines on most tasks, especially on Edge Number Query, where results vary greatly with complexity. Besides, some LVLMs achieve peak performance at intermediate complexity levels on some tasks, but struggle with more complex graphs. This reflects a balance between information richness and complexity of visual graphs, whereas higher complexity likely overwhelm the LVLMs’ reasoning capacity, attention mechanisms, or ability to generalize due to training biases or compounded errors in large graphs. In addition, different complexity dimensions have varying effects on the LVLMs performance on different tasks. For instance, on Nodes Number Query, LVLMs are minimally affected by the number of edges while strongly affected by the number of nodes and average degree. More results are shown in Figures 18-23 in the Appendix due to space limit, and the overall trend remains similar to the above findings.

### v2

@@ Abstract
Large Vision-Language Models (LVLMs) have demonstrated remarkable performance across diverse tasks. Despite great success, recent studies show that LVLMs encounter substantial limitations when engaging with visual graphs. To study the reason behind these limitations, we propose VGCure, a comprehensive benchmark covering 22 tasks for examining the fundamental graph understanding and reasoning capacities of LVLMs. Extensive evaluations conducted on 14 LVLMs reveal that LVLMs are weak in basic graph understanding and reasoning tasks, particularly those concerning relational or structurally complex information. Based on this observation, we propose a structure-aware fine-tuning framework to enhance LVLMs with structure learning abilities through three self-supervised learning tasks. Experiments validate the effectiveness of our method in improving LVLMs’ performance on fundamental and downstream graph learning tasks, as well as enhancing their robustness against complex visual graphs. [^1] Our dataset and code are available at https://github.com/AAAndy-Zhu/VGCure.
@@
Yingjie Zhu1,2, Xuefeng Bai1, Kehai Chen1, Yang Xiang2, Jun Yu1, Min Zhang1 1Harbin Institute of Technology, Shenzhen, China 2Peng Cheng Laboratory, Shenzhen, China zhuyj@stu.hit.edu.cn, {chenkehai,baixuefeng,zhangmin2021}@hit.edu.cn xiangy@pcl.ac.cn, yujun@hdu.edu.cn
@@ 1 Introduction
Graphs serve as a fundamental data structure across a wide range of domains, including social network analysis Schweimer et al. (2022), recommendation systems Zhang et al. (2023), knowledge graphs Zhang et al. (2024b), chemistry Cao et al. (2024) and semantic reasoning (Bai et al., 2022). Existing methods have achieved great success in enhancing understanding and reasoning abilities in graph-based tasks Kim et al. (2023a); Chen et al. (2024b). However, these approaches typically focus on specific graph types or tasks, posing challenges in designing versatile systems that are suitable for various tasks and graphs across diverse domains.
@@
Recently, Large Vision-Language Models (LVLMs) have exhibited outstanding performance across a wide range of downstream tasks by unifying various inputs in the form of images and processing them with human-like understanding and reasoning abilities Zhu et al. (2024); Zheng et al. (2024); Chen et al. (2024a). This triggers a growing interest in employing LVLMs for graph learning problems, as the vision modality offers a natural and intuitive way for comprehending structural information and facilitating general graph reasoning Poklukar et al. (2022). However, recent studies Wei et al. (2024); Li et al. (2024c); Ai et al. (2024) reveal significant challenges for LVLMs in graph-based learning tasks, where LVLMs achieve less than 15% accuracy on mathematical graph reasoning tasks, markedly below their performance in image and text reasoning Li et al. (2024c). Therefore, a natural challenge arises: why do LVLMs fail in graph learning, and how to enhance LVLMs to process graphs like professionals?
@@
To address this challenge, we begin by identifying a research gap in existing evaluations that limit our understanding of LVLMs. Current work primarily focuses on graph theory problems or multi-hop reasoning tasks Wei et al. (2024); Ai et al. (2024), which are complex and require a combination of diverse cognitive abilities. Thus it remains uncertain whether LVLMs have strong fundamental understanding and reasoning capabilities in processing visual graphs—such as recognizing the basic components of a graph and making basic logical inferences based on graph data—which is crucial for determining whether the limitations stem from a lack of fundamental abilities or other higher-order capabilities.
@@
To this end, we present the Vision Graph Comprehensive Understanding and REasoning benchmark, VGCure, designed to thoroughly evaluate the fundamental understanding and reasoning capabilities of LVLMs on visual graphs. As shown in Fig.1, VGCure evaluates the fundamental capabilities of LVLMs diverse challenges, including 9 graph understanding tasks and 13 graph reasoning tasks. Moreover, VGCure features 10 anonymized graph structures from both synthetic and real-world sources, offering a robust testbed for assessing LVLMs’ proficiency in handling diverse graphs. Experiments on 14 representative LVLMs show that current LVLMs exhibit weak fundamental understanding and reasoning capabilities on visual graphs, especially in capturing relational information and dealing with structurally complex visual graphs.
@@
Motivated by the above observations, we further introduce MCDGraph, a novel structure-aware fine-tuning framework to enhance structure learning capabilities of LVLMs through three self-supervised learning tasks: 1) masked graph infilling, 2) contrastive graph discrimination, and 3) graph description. Experiments show that MCDGraph significantly improves LVLMs’ graph understanding and reasoning abilities, especially on edge-related tasks and graph-related downstream tasks. Further analysis demonstrates that our method also enhances LVLMs’ robustness and generalization to visual graphs with complex structure and unseen styles. The contributions of this work can be summarized as follows:
@@
  • Through extensive experiments on 14 LVLMs, we reveal LVLMs’ limitations in basic graph understanding and reasoning, especially for tasks concerning relational or structural information.
  • We propose a self-supervised framework to enhance LVLMs’ ability to capture structural information in visual graphs. Experiments validate its effectiveness on both fundamental and downstream graph learning tasks.
@@
Table 2: Examples for each task in VGCure. These samples all correspond to the graph shown in Fig.LABEL:fig:example_graphs(d).
@@ 2 The VGCure Benchmark
To evaluate LVLMs’ fundamental graph understanding and reasoning capabilities, we introduce VGCure, a large-scale multimodal graph benchmark with 22 challenging tasks. VGCure features 10 graph types, both synthetic and real-world, to assess LVLMs’ performance on diverse graphs. The graphs are anonymized to minimize the impact of pre-existing LLM knowledge on core reasoning abilities, promoting knowledge-free reasoning (Hu et al., 2024). Tab.1 compares VGCure with three recent benchmarks. It is evident that VGCure excels in fundamental graph understanding and reasoning capabilities. Furthermore, VGCure offers a comprehensive evaluation through more varied graph types, tasks, and test samples.
@@ 2.1 Graph Structure Generation
We begin by collecting a wide variety of graph structures for generating visual graphs and challenging tasks. Following Fatemi et al. (2024), we first employ NetworkX (Hagberg et al., 2008) to generate a diverse set of random synthetic structures, including Erdős-Rényi (ER) graphs (Erdős and Rényi, 1959), scale-free networks (SFN) (Barabási and Albert, 1999), Barabási–Albert (BA) model (Albert and Barabási, 2002), stochastic block model (SBM) (Holland et al., 1983) and star graphs. In addition, we extract anonymized structures from real-world graphs in GraphArena (Tang et al., 2024), including DBLP, Social Network, DBpedia, Openflights and PubChemQC. All the entity and relation names in each graph are replaced with generic names to eliminate the impact of the model’s internal knowledge on reasoning. After initializing the graph structure, we use the Graphviz (Ellson et al., 2002) to generate concise directed and undirected visual graphs.
@@ 2.2 Tasks Design
To thoroughly assess the abilities of LVLMs in fundamental graph structure understanding and reasoning, the proposed VGCure encompasses the following categories of tasks. Tab.2 presents examples for each task.
Graph Understanding: The graph understanding tasks involve analyzing and extracting structural, relational, and property-based information from the visual graph, which aims to gain insights into the composition and topology of the graph, including its nodes, edges, connectivity, and the relationships among its components.
@@
  • Degree Query (DQ): Calculate the degree of the specified node.
@@
Graph Reasoning: The reasoning tasks focus on exploring the knowledge-free reasoning ability of LVLMs on visual graphs. To differentiate from graph understanding tasks, we designed a series of 2-hop reasoning tasks.
@@
  • Conjunctive Relation Query (CR): Retrieve the entities satisfying both two independent relationship constraints with two distinct entities.
@@
For each task, we construct one QA sample and two fact checking (FC) samples (with labels of True and False, respectively) automatically based on the template, except for CN and DC which have only fact checking samples due to the strong similarity between the two samples. The total number of final samples is 223,646. More details about the design logic and statistics of VGCure can be found in Appendix A and Tab.7.
@@ 3.1 Experimental Setup
We conduct evaluation on 13 open-source LVLMs, including InternLM-XComposer2.5-7B (Zhang et al., 2024a), InternVL2-8B, Llama3.2-11B-Vision-Instruct, LLaVA1.5-7B (Liu et al., 2024), LLaVA1.5-13B (Liu et al., 2024), LLaVA-NeXT-7B (Li et al., 2024b), LLaVA-OneVision-7B (Li et al., 2024a), MiniGPT-v2 (Chen et al., 2023), Monkey (Li et al., 2024d), mPLUG-Owl3-7B (Ye et al., 2024), Qwen-VL (Bai et al., 2023), Qwen2-VL-7B-Instruct (Wang et al., 2024) and SPHINX (Lin et al., 2023). Meanwhile, we also evaluate the performance of the GPT-4o-mini, which is a strong closed-source LVLM [^2] We excluded GPT-4o as a baseline due to its high cost.. Due to the high cost of GPT-4o-mini, we randomly select 50 graphs from each graph structure for testing. For all methods, the zero-shot setting is adopted during evaluation. More details can be found in Appendix B.
Figure 2: Model performance (F1/Acc) on representative QA samples across various graph structures and tasks , where OF, PCQC and SN denotes Openflights, PubChemQC and Social network, respectively.
@@ 3.2 Main Result
Graph Understanding Tab.3 and Tab.4 present the evaluation results for question answering (QA) and fact checking (FC), respectively. We report the averaged results across various graph structures. In general, most LVLMs struggle to precisely understand the structural and relational information in visual graphs. In specific, (I) Among the graph understanding tasks, Node Number Query (NNu) and Directedness Check (DC) are the easiest for most LVLMs. This indicates that most LVLMs can accurately capture the number of nodes and directedness information within the visual graph. (II) All LVLMs struggle with Edge Number Query (EN) and Neighbor Query (NQ) tasks, with the highest accuracy of 16.38% and F1 score of 30.81% , respectively. This indicates that current LVLMs are weak in understanding relational and structural information. (III) Even with a similar number of parameters, the graph understanding abilities of open-source LVLMs vary significantly, with Qwen2-VL and InternVL2 showing better performance in both QA and FC samples. (IV) For the same task, LVLMs perform differently on QA and FC samples, likely due to different ways in reasoning and understanding required by each task (Thorne et al., 2018). (V) The closed-source LVLM, GPT-4o-mini, offers no significant advantages and even underperforms open-source LVLMs, especially on FC tasks.
Graph Reasoning Based on graph reasoning results in Tabs.3 and 4, it can be observed that, (I) Compared to graph understanding, the graph reasoning tasks are more challenging and the overall performance of LVLMs is worse on both QA and FC samples. This may be a knock-on effect due to deficiencies in visual graph understanding. (II) Among all open-source LVLMs, InternVL2 and Llama3.2-Vision achieve better performance and LLaVA1.5-7B perform the worst on graph reasoning tasks. (III) All the LVLMs perform poorly on Nested Relation Query (NR) for both QA and FC samples, which is similar to the observation in the graph understanding task. This suggests that LVLMs are deficient in recognizing edges and understanding structural information within visual graphs. (IV) On the QA samples, the performance of different LVLMs on Connective Path Query (CP) varies widely. SPHINX, Monkey and Qwen-VL demonstrate almost no ability to recognize paths between two specific nodes in the visual graph. (V) Similarly, GPT-4o-mini underperforms in most tasks compared to open-source LVLMs, except for the Connective Path Query (CP) task.
@@ 3.3 Impact of Structures
Inspired by Fatemi et al. (2024), we explore how graph structure affects LVLMs’ ability to understand and reason on visual graphs. Fig.2 compares the performance of the top five LVLMs on QA samples across various structures. Obviously, the graph structure significantly impacts LVLMs’ performance on most tasks. Notably, all LVLMs perform well on PCQC, which has a simpler structure with fewer nodes and edges (averaging 5.45 nodes and 4.76 edges), and weaker on BA, which has the highest edge count in VGCure (averaging 21.02 edges). Furthermore, on different tasks, the performance of LVLMs is affected differently by the graph structure. EN and CP show larger performance variations across graph structures, whereas NNu and NN show smaller differences. More results are shown in Figs.23 and 24, the overall trends are similar to above findings.
@@ 3.4 Impact of Complexity
We further discuss the impact of graph complexity on the LVLMs’ ability to understand and reason over the visual graph by considering the number of edges, number of nodes, and average degree. Fig.LABEL:fig:complexity_results shows a performance comparison of five LVLMs on QA samples across representative tasks and complexity levels. As complexity increases, LVLMs’ performance declines, especially on EN, where results vary significantly. Besides, some LVLMs perform best at intermediate complexity, but struggle with more complex graphs. This reflects a balance between information richness and complexity of visual graphs, whereas higher complexity likely overwhelm the LVLMs’ abilities to reason or generalize due to the complex information within large graphs. In addition, different complexity dimensions affect LVLMs’ performance on various tasks differently, such as NNu, where LVLMs are more influenced by the number of nodes and average degree than by edges. More results in Figs.25-30 show the similar trend.
3.5 Why LVLMs Fail on Fundamental Tasks?
To explore in depth whether the failure of LVLMs on the fundamental task is due to their weak perception or other higher-order ability, we compare the performance of LVLMs and the corresponding backbone LLMs that take the textual graph as input. As Fig.LABEL:tab:text_results shows, despite the disparity in the performance of the individual models, LVLMs perform weaker than the corresponding backbone LLMs on both understanding and reasoning tasks. In particular, LVLMs’ performance degrades the most on Edge Number Query (EN), which also illustrates the lack of ability of LVLMs to capture structural information in the visual graph. This indicates that the failure of LVLMs is partly attributed to their weaker visual graph perception. In addition, the performance of the backbone LLMs on these tasks remains undesired, with all accuracies below 45%, this suggests that the backbone LLMs exhibit limited capabilities in understanding and reasoning on graph data. More details and results are presented in Appendix C.1.
In addition, by analyzing the generation results of LVLMs, we find that their main errors can be grouped into the following six categories: Relation Misunderstanding, Complexity Limitation, Structural Hallucination, Question-based Answering, Off-target Generation, and Format Error. Due to space constraint, we provide the detailed description in Appendix C.2.